When 0.540
g of sodium metal is added to an excess of hydrochloric acid, 5610
J of heat are produced. What is the enthalpy of the reaction as written?

2Na(s)+2HCl(aq)⟶2NaCl(aq)+H2(g)

Enthalpy of reaction
1
Expert's answer
2020-07-13T07:30:01-0400

"The\ enthlapy\ of\ this\ reaction\ is \ the\ released\ energy\ thus\ \Delta H \ is \ neagive\ sign.\newline 2Na(s)\ \ +\ 2HCl(aq)\ \to \ 2NaCl(aq)\ \ +\ H2(g)\ \ +\Delta??\\ \\\"

Number of moles of sodium "=\dfrac{0.540.g}{22.99.g.mol^{-1}}"

"=0.0235.\ mol"

Therefore,

"\Delta H^0rxn=\dfrac{5610.J}{0.0235.mol} *\dfrac{1}{1000}" since enthalpy per mole is expressed in "kJ\/mol"


"=-238.7.kJ. mol"